# Binaural Target Speaker Extraction using HRTFs

### v2

@@ Abstract
Index Terms:
Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
I Introduction
@@ III-A Architecture
To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor is passed through four consecutive self-attention layers and a fully-connected layer, preparing it for the decoding stage. A block diagram of the proposed architecture is presented in Fig. 1.
@@
Fig. 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated binaural desired signal.
III-B Loss function
We initially trained the proposed model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
@@ IV-A Data Generation
The RIEC HRTF Dataset [19, 20]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA.
We employed the SofaMyRoom framework [23] to simulate both anechoic and reverberant conditions, following the formulation in (6). This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time. It also supports zero-order simulations for anechoic scenarios. The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room. The overall simulation setup is illustrated in Fig. 2.
@@ IV-C Metrics
Table I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples. Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap . Best scores are highlighted in bold.
@@
IV-D Competing Methods
IV-E Results
In Table I, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions. In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE. Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
Remarkably, both DNN-based methods demonstrate strong cross-dataset generalizability. Although the models were trained exclusively on WSJ0-CSR1, they perform well on other corpora, including a French dataset. We attribute this generalization to the absence of speaker-specific priors and to the reliance on a known HRTF database.
@@ IV-E Results
Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI. The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation. By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation.
Figure 3 shows the joint ITDILD p.d.f.s across methods. The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues. In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations. The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f. becomes more concentrated, indicating dominance of the direct path in the output of the proposed method. Overall, it is evident that the proposed approach achieves this more effectively than competing techniques. Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig. 3. Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec. IV-C. The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
@@
Finally, to quantify the spatial resolution of our model, we employ a “scanning” procedure: signals are iteratively extracted using HRTFs at azimuths within $90^{\circ},90^{\circ}]$ (with fixed elevation), and the SI-SDR of each extraction is computed against the target. Plotting these values produces a beam-pattern-like diagram (Fig. 4) that exhibits a clear peak at the true DOA. The width of the main lobe further characterizes the model’s robustness when the HRTF cue deviates slightly from the ground-truth.
V Conclusions

### v3

@@ Abstract
Index Terms—  Binaural target speaker extraction; Head-related transfer function; Complex-valued Neural network.
1 Introduction
@@ 3.1 Architecture
To capture cross-channel (left and right) information in the HRTF, we apply a self-attention mechanism that enables the channels to attend to each other through learned attention weights. The resulting matrix is flattened into a vector and passed through a fully connected layer to match the dimension of the encoded mixed signal. The output of this process serves as the HRTF’s embedding. Similarly, self-attention is applied to the encoded mixed signal. We apply element-wise multiplication between the (replicated) HRTF’s embedding and each of the time frames of the encoded mixed signal, thus effectively imposing the clue over the mixed signal. Finally, the resulting tensor undergoes four consecutive self-attention layers and a fully connected layer, preparing it for the decoding stage. A block diagram of the proposed architecture appears in Fig. 1.
@@
Fig. 1: A block diagram of the proposed method, where $\bm{x}_{b}$ represents the mixed binaural signal in the STFT domain, $\bm{h}_{\text{hrtf}}{(\theta_{d},\phi_{d},k)}$ denotes the (frequency-domain) HRTF of the desired speaker’s DOA, and $\hat{\tilde{\bm{s}}}_{d}$ represents the STFT of the estimated desired signal.
3.2 Loss function
We initially trained the model using a single loss function, the Scale Invariant Signal to Distortion Ratio (SI-SDR) loss, as in (3).
@@ 4.1 Data Generation
The RIEC HRTF Dataset [19, 20]: We utilized the RIEC database, which comprises measurements from 105 subjects, each with 865 directions. For this study, we focused on a single subject. The Spatially Oriented Format for Acoustics (SOFA) HRTF Toolbox [21, 22] facilitates easy retrieval of the appropriate HRTF for any desired DOA. We employed the SofaMyRoom framework [23] to simulate both anechoic and reverberant conditions, following the formulation in (6). This framework positions a listener, characterized by their measured HRTF, inside a shoebox-shaped room with configurable wall materials to achieve a specified reverberation time. It also supports zero-order simulations for anechoic scenarios. The reverberation time was sampled as $T_{60}\sim\mathcal{U}[0.2,0.8]$, and both speaker and listener positions were drawn uniformly within the room. The overall simulation setup is illustrated in Fig. 2.
@@ 4.3 Metrics
Table 1: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples. Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap . Best scores are highlighted in bold.
@@
s
@@
4.4 Competing Methods
4.5 Results
In Table 1, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions. In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE. Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
Both DNN-based methods show strong cross-dataset generalization: though trained only on WSJ0-CSR1, they perform well on other corpora, including French, as shown in Table 1. We attribute this to the absence of speaker-specific priors and the reliance on a known HRTF database.
Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI. The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation. By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation. Figure 3 shows the joint ITDILD p.d.f.s across methods. The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues. In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations. The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f. becomes more concentrated, indicating dominance of the direct path in the output of the proposed method. Overall, it is evident that the proposed approach achieves this more effectively than competing techniques. Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig. 3. Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec. 4.3. The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
@@ 4.5 Results
Finally, to quantify spatial resolution, we use a “scanning” procedure: signals are extracted with HRTFs at azimuths in $90^{\circ},90^{\circ}]$, and the SI-SDR of each is computed against the target. Plotting these values yields a beam-pattern-like diagram (Fig. 4) with a clear peak at the true DOA. The main lobe width characterizes the model’s stability when the HRTF clue deviates mildly from the ground truth.
5 Conclusions